OpenStudy (anonymous):

9. Which of the following is the factorization of the quadratic below? x2 - 5x + 4 A. (x - 2)(x - 2) B. (x + 4)(x + 1) C. (x - 4)(x - 1) D. (x - 4)(x + 1)

ganeshie8 (ganeshie8):

pick two numbers such that they : 1) multiply to 4 2) add up to -5

ganeshie8 (ganeshie8):

knw how to pick such two numbers ?

OpenStudy (anonymous):

nope @ganeshie8

ganeshie8 (ganeshie8):

try picking them from factors of 4 : 4*1 = 4 however 4+1 = 5 which is not negative 4 so this wont work

ganeshie8 (ganeshie8):

next try -4 and -1 (-4)(-1) = 4 (-4) + (-1) = -5 this seem to fit the requirement, eh

ganeshie8 (ganeshie8):

\[x^2-5x+4 = (x-4)(x-1)\]

OpenStudy (anonymous):

then just solve?

ganeshie8 (ganeshie8):

then you're done look at ur options

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

Here is how you can factor many quadratics with ease.

OpenStudy (anonymous):

was i right though dude?

OpenStudy (anonymous):

B is not correct no.

OpenStudy (anonymous):

frick it was a wasnt it

OpenStudy (anonymous):

You can check it by using F.O.I.L. and seeing what it comes out too. B. (x+4)(x+1) so x^2 +5x + 4 is what we get.

OpenStudy (anonymous):

F.O.I.L. stands for First. Outside. Inside. Last.